Plastic parts produced by an injection-molding operation are checked for conformance to specifications. Each tool contains 11 cavities in which parts are produced, and these parts fall into a conveyor when the press opens. An inspector chooses 1 part(s) from among the 11 at random. Three cavities are affected by a temperature malfunction that results in parts that do not conform to specifications.

a) What is the probability that the inspector finds exactly one nonconforming part?

b) What is the probability that the inspector finds at least one nonconforming part?
Answer #1

(a) There are 11C1 = 11 ways for the inspector to pick 1 part to check.
When exactly one is bad: there are 3 ways to pick the bad part

So 3/11 is the required probability

b) the probability that the inspector checks only good parts is 8C1/11C1 = 8/11
therefore the probability that he finds at least one bad is 1 - p(all good parts)= 1-8/11 = 3/11